what is logx + log (x-0.21) = -2?
Answer by stanbon(57347)

what is logx + log (x-0.21) = -2?
log[x(x-0.21)) = -2
------------
x(x-0.21) = 10^-2
x^2 - 0.21x - 0.01 = 0
(x + 0.04)(x - 0.25) = 0
Positive solution:
x = 0.25
=================
Cheers,
Stan H.